You are trying to decide if you have enough money to retire. You are expecting $1600 a month from Social Security. You have $1 million in your retirement savings, and you only want to use the earnings to live on, which is at 3.5% annual rate. You are currently making $64,000 a year and plan to live 80% of that amount. Do you have enough to retire based on these numbers?
Found 2 solutions by jorel555, addingup:
Answer by jorel555(1290)

64000 x .8=51200
1000000 x 3.5%=35000
1600 x 12=19200
35000+19200=$54200 annually to live on; which exceeds your goal of $51200. ☺☺☺☺

Answer by addingup(3677)

You plan to live on:. . . . . . . . . . . . . . . . . .64,000*0.8 = 51,200
:
soc sec . . . . . . . . . . 1,600*12 = 19,200
1 mill at 3.5% . . . . . . . . . . . . = 35,000
. . . . . . . . . . . . . . . . . . . . . . . .--------
total income . . . . . . . . . . . . . = 54,200
:
You have 54,200-51,200 = 3,000 per year to either save or go on a nice vacation or... squander gambling if that's what makes you happy :-)